OpenStudy (anonymous):

15. Factor completely: 4x^2 + 13x + 3 A. (2x + 3)(2x + 1) B. (4x + 1)(x + 3) C. (2x + 3)(2x + 3) D. (4x + 3)(x + 1)

OpenStudy (tkhunny):

Have you considered multiplying the two factors and seeing what you get?

OpenStudy (anonymous):

4x^2+13x+3 =4x(x+3)+(x+3) =(4x+1)(x+3)=B